# Zero-shot denoising via neural compression: Theoretical and algorithmic framework

### v1

@@ 1 Introduction
Denoising is a fundamental problem in classical signal processing and has recently gained renewed attention from the machine learning community. Let ${\bm{x}}=(x_{1},\ldots,x_{n})\in\operatorname{\mathbb{R}}_{+}^{n}$ denote a non-negative signal of length $n$, where signal ${\bm{x}}$ is not observable in many systems. Instead, we observe a noisy version ${\bm{y}}=(y_{1},\ldots,y_{n})$, where the observations are conditionally independent given ${\bm{x}}$, and each entry is distributed according to a common conditional distribution:
@@
We assume that the noise mechanism is memoryless (independent across coordinates) and homogeneous (identical across entries). The goal of a denoising algorithm is to estimate ${\bm{x}}$ from the noisy observations ${\bm{y}}$. Given its prevalence in imaging and data acquisition systems, denoising has been a central topic in signal processing for decades. Classical denoising methods rely on explicit structural assumptions about the underlying signal ${\bm{x}}$, often hand-crafted by domain experts [1, 2, 3, 4, 5, 6, 7, 8]. In contrast, recent advances in machine learning have enabled a new class of data-driven denoising algorithms. These methods learn the optimal denoising function from data, leveraging statistical patterns directly from signal and noise distributions.
@@
To mitigate the reliance on paired clean and noisy samples, several self-supervised denoising methods have been developed that learn directly from noisy observations, without access to clean ground truth signals [9, 10, 11, 12, 13]. While these approaches alleviate the supervision requirement, they typically depend on access to large collections of noisy data and often yield suboptimal performance compared to methods trained with clean targets. Moreover, the absence of clean supervision necessitates the use of complex neural architectures and training schemes, which can make these methods computationally demanding and difficult to optimize in practice.
@@
In lossy compression, the goal is to represent signals from a target class using discrete encodings with minimal distortion. When applied to noisy data, the intuition is that a lossy compressor—operating at a distortion level matched to the noise—will favor reconstructions close to the original clean signal. While this approach has a strong theoretical foundation [14, 15], classical compression-based denoisers have shown limited empirical success, particularly for natural image denoising.
In this work, we revisit this idea in light of recent progress in neural compression, where learned encoders and decoders have demonstrated strong rate-distortion performance across a variety of image domains [16, 17]. Building on this foundation, we propose a zero-shot denoising method that we call the Zero-Shot Neural Compression Denoiser (ZS-NCD). Unlike traditional neural compression models that are trained on large corpora of clean high-resolution images, ZS-NCD learns directly from a single noisy input image. Specifically, we extract overlapping patches from the noisy image, and train a neural compression network on those patches alone—without any clean supervision or prior dataset. Once trained, the denoiser is applied to all patches from the same image, and the final output is obtained by averaging the predictions in overlapping regions. This approach is illustrated in Figure 1.
@@
Figure 2: Zero-shot denoising of Kodim05 with AWGN ($\sigma=25$). Left: PSNR versus training iterations for zero-shot denoisers. Performance of BM3D [18] and Restormer [19] are included as a classical baseline and as a supervised empirical upper bound, respectively. Right: Visual reconstructions with PSNR/SSIM above each image. Compression-based denoising based on JPEG-2K [20] achieves inferior performance. Learning-based zero-shot denoisers often struggle with either overfitting or high bias. DIP [21] and DD [22] require early stopping to avoid overfitting. ZS-N2S [12] and S2S [23] struggle with high-resolution color images, and ZS-N2N [24] often produces noisy outputs with potential overfitting. BM3D tends to oversmooth the denoised image. In contrast, ZS-NCD avoids these issues.
@@ 2 Related work
Supervised learning-based denoisers such as DnCNN [25] and Restormer [19] achieve state-of-the-art performance across various noise models, but require large datasets of paired clean and noisy images—often impractical in real-world settings. To avoid clean images, self-supervised methods have been proposed, including Noise2Noise [10], Noise2Self [12], Noise2Void [11], Noise2Same[26] and Noise2Score [13], which only use noisy images for training. However, their reliance on large noisy datasets remains a limitation. Zero-shot denoisers address this by training on a single noisy image. These include (i) untrained networks like DIP [27] and Deep Decoder [22], and (ii) single-image adaptations of self-supervised methods, e.g., ZS-N2N [24], ZS-N2S [12] and its augmented variant with ensembling, S2S [23]. DIP-based models avoid masking and leverage full-image context, but require early stopping or under-parameterization to avoid overfitting. Self-supervised variants suffer from masking-induced information loss. Hybrid approaches, such as masked pretraining-based method [28], uses external datasets for training and perform zero-shot inference, thus falling outside the zero-shot setting studied here.
@@
Learning-based lossy compression, often referred to as neural compression, uses an autoencoder architecture combined with an entropy model to estimate and constrain the bitrate at the bottleneck [29, 30]. These methods have significantly outperformed traditional codecs, particularly in image [29, 31, 32, 33, 34] and video [35, 36, 37] compression. In addition to these standard settings, several works have also explored applying neural compression to noisy data, either by adapting neural compression models for more efficient encoding of noisy images [38, 39, 40, 41].
@@
Compression-based denoising leverages the insight that structured signals are inherently more compressible than their noisy counterparts. This connection was formalized by Donoho [14], who introduced the minimum Kolmogorov complexity estimator, and further refined by Weissman et al. [15], showing that, under certain conditions on both the signal and the noise, optimal lossy compression of a noisy signal-followed by suitable post-processing-can asymptotically achieve optimal denoising performance. Prior to these theoretical developments, early empirical methods such as wavelet-based schemes [42, 43, 44] and MDL-inspired heuristics [45] had explored this principle. Nevertheless, traditional compression-based denoisers have generally underperformed in high-dimensional settings such as image denoising.
Learning-based joint compression and denoising using neural compression has been explored in recent works [46, 47], where the goal is to achieve lower rate in compression. The empirical application of training neural compression for AWGN denoising was also proposed in [48]. Training the neural compression models in these works requires a dataset of images. In contrast, our proposed ZS-NCD is a two-step denoiser based on neural compression, trained on a single noisy image. It achieves state-of-the-art performance across both AWGN and Poisson noise models. Moreover, we contribute new theoretical results that advance the foundations of compression-based denoising.
@@ 3 Compression-based denoising: Theoretical foundations
Let $\mathcal{Q}\subset\operatorname{\mathbb{R}}^{n}$ denote the signal class of interest, such as vectorized natural images of a fixed size. A lossy compression code for $\mathcal{Q}$ is defined by an encoder-decoder pair $(f,g)$, $f:\mathcal{Q}\to\{1,\ldots,2^{R}\}$, and $g:\{1,\ldots,2^{R}\}\to\operatorname{\mathbb{R}}^{n}$. The performance of a lossy code is characterized by: i) Rate $R$, indicating the number of distinct codewords; ii) Distortion $\delta$, defined as the worst-case per-symbol mean squared error (MSE) over the signal class:
@@
$\mathcal{C}=\{g(i):i=1,\ldots,2^{R}\}\subset\operatorname{\mathbb{R}}^{n}.$
@@
We begin by analyzing the performance of compression-based ML denoising under AWGN. The following result provides a non-asymptotic upper bound on the reconstruction error in terms of the compression rate and distortion. All proofs can be found in Appendix 6.
@@
${1\over\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq\sqrt{\delta}+2\sigma_{z}\sqrt{(2\ln 2)R\over n}(1+2\sqrt{\eta}),$ (4)
@@
Let $\mathcal{Q}_{n}$ denote the set of $k$-sparse vectors in $\operatorname{\mathbb{R}}^{n}$ satisfying $\|{\bm{x}}\|_{2}\leq 1$. Fix a parameter $\eta\in(0,1)$, and suppose ${\bm{y}}={\bm{x}}+{\bm{z}}$ where ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}I_{n})$. Then, there exists a family of compression codes such that, when used with the denoiser defined in (1), the estimate $\hat{\bm{x}}$ satisfies
$\frac{1}{n\sigma_{z}^{2}}\|\hat{{\bm{x}}}-{\bm{x}}\|_{2}^{2}\leq C\cdot\frac{k\log_{2}n}{n}+\gamma_{n},$
with probability at least $1-\frac{4}{(kn^{3k/2})^{\eta}}$. Here, $\gamma_{n}=o(1)$ and $C=4\ln 2(1+2\sqrt{\eta})^{2}$.
To contextualize this result, consider the asymptotic setting where $n\to\infty$ and $\sigma_{z}^{2}\to 0$. For an i.i.d. Bernoulli–Gaussian source with ($\mathbb{P}\left\lparen x_{i}\neq 0\right\rparen=p$, the minimum mean squared error (MMSE), normalized by the noise power $\sigma_{z}^{2}$, converges to $p$ [49]. In contrast, our result is non-asymptotic and worst-case: it shows that the MSE of compression-based denoising is bounded by $C\cdot\frac{k\log n}{n}$. In the probabilistic setting where $k/n\approx p$, the two results exhibit consistent scaling with sparsity.
@@
Consider the same setup of lossy compression as in Theorem 1. Assume that for any ${\bm{x}}\in\mathcal{Q}$, $x_{i}\in(x_{\min},x_{\max})$, where $10<x_{\min}<x_{\max}<$ 1. Assume that $y_{1},\ldots,y_{n}$ are independent with $y_{i}\sim\mathrm{Poisson}(\alpha x_{i})$. Let $\hat{\bm{x}}$ denote the solution of (2). Let $C_{1}={x_{\max}^{5}/(x_{\min}^{2})}$ and $C_{2}={x_{\max}^{2}\over x_{\min}^{3}}\beta\sqrt{({4\over\ln 2})}(\sqrt{1+\eta}+\sqrt{\eta})$. Then, with a probability larger than $1-2^{-\eta R+2}$,
@@
Consider the same setup as in Theorem 2. Let $\hat{\bm{x}}$ denote the solution of (3). Let $C=\left\lparen 4\sqrt{\ln 2}\right\rparen\left\lparen\sqrt{1+\eta}+\sqrt{\eta}+1\right\rparen$ ). Then, with a probability larger than $1-2^{-\eta R+2}$,
${1\over n}\|{\bm{x}}-\hat{\bm{x}}\|_{2}^{2}$ $\leq\delta+C\sqrt{R\over n\alpha}$ (6)
@@
Theorems 2 and  3 show that, in the case of Poisson noise, there is no loss in performance, when the more computationally efficient MSE loss function is minimized instead of the ML loss function. This result is also consistent with our simulations reported later in Section 5.
@@ 4 Zero-shot compression-based denoiser
Let $\mathcal{P}_{(i,j)}:\operatorname{\mathbb{R}}^{h\times w}\to\operatorname{\mathbb{R}}^{k\times k}$ denote the patch extraction operator, which returns a $k\times k$ patch whose top-left corner is at pixel $(i,j)$. Let $f_{\bm{\theta}_{1}}$ and $g_{\bm{\theta}_{2}}$ denote the encoder and decoder networks, parameterized by weights $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, respectively. Define $\mathcal{I}$ as the set of all coordinates $(i,j)\in\{1,\ldots,h-k+1\}\times\{1,\ldots,w-k+1\}$ from which a valid $k\times k$ patch can be extracted.
@@
$(\hat{\bm{\theta}}_{1},\hat{\bm{\theta}}_{2})=\operatorname*{arg\,min}_{(\bm{\theta}_{1},\bm{\theta}_{2})}\sum_{(i,j)\in\mathcal{I}}\left(\mathcal{L}_{K}(g_{\bm{\theta}_{2}}(f_{\bm{\theta}_{1}}(\mathcal{P}_{(i,j)}({\bm{y}}))),\mathcal{P}_{(i,j)}({\bm{y}}))-\lambda\log\mathbb{P}\left\lparen f_{\bm{\theta}_{1}}(\mathcal{P}_{(i,j)}({\bm{y}}))\right\rparen\right),$ (7)
where ($\mathbb{P}\left\lparen f_{\bm{\theta}_{1}}(\mathcal{P}_{(i,j)}({\bm{y}}))\right\rparen$ ) denotes the likelihood (or entropy model) of the latent code produced by the encoder, and $\lambda>0$ is a hyperparameter controlling the trade-off between fidelity and compressibility. In (7), $K=k^{2}$, and the function $\mathcal{L}_{K}:\operatorname{\mathbb{R}}^{K}\times\operatorname{\mathbb{R}}^{K}\to\operatorname{\mathbb{R}}_{+}$ is a distortion loss determined by the noise model, as defined in Section 3. For example, in the AWGN case, $\mathcal{L}_{K}$ corresponds to the squared $\ell_{2}$ norm of the distance between a noisy patch and its neural compression reconstruction. Note that $f_{\bm{\theta}_{1}}$ maps the input into a discrete latent space, which is non-differentiable and thus incompatible with standard gradient-based optimization. To address this, we follow the neural compression framework of [29], using a continuous relaxation during training (e.g., uniform noise injection) and applying actual discretization only at test time. The entropy term $\mathbb{P}$ is modeled using a factorized, non-parametric density [31].
@@
1:  Initialize: $\lambda^{(0)}$
2:  for iteration $k=0,1,\dotsc,K_{\max}-1$ do
3:     Estimate $\hat{\bm{x}}^{(k)}$ from (8), and update $\lambda$:
$\lambda^{(k+1)}:=\begin{cases}2\lambda^{(k)},&\text{if }\|\hat{\bm{x}}^{(k)}-\bm{y}\|^{2}>n\sigma_{z}^{2}\\ \lambda^{(k)}/2,&\text{otherwise}\end{cases}$
4:  end for
5:  Return: $\lambda^{*}$
@@
Setting the hyperparameter $\lambda$
@@
The ZS-NCD objective in (7) includes a hyperparameter, $\lambda$, which balances reconstruction fidelity and compressibility. Interpreted through the lens of lossy compression, varying $\lambda$ allows the model to explore different rate-distortion trade-offs. However, in the context of denoising, our goal is not compression but accurate signal recovery from the noisy observation ${\bm{y}}$. This raises the central question: how should $\lambda$ be selected to optimize denoising performance?
@@
Let $\hat{\bm{x}}$ denote the output of the ZS-NCD denoiser, and consider the AWGN model ${\bm{y}}={\bm{x}}+{\bm{z}}$ with ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}I_{n})$. Then,
$\textstyle{1\over n}\mathbb{E}\left[\|{\bm{y}}-\hat{\bm{x}}\|_{2}^{2}\right]={1\over n}\mathbb{E}\left[\|{\bm{y}}-{\bm{x}}+{\bm{x}}-\hat{\bm{x}}\|_{2}^{2}\right]=\sigma_{z}^{2}+{1\over n}\|{\bm{x}}-\hat{\bm{x}}\|_{2}^{2}+{2\over n}\mathbb{E}\left[{\bm{z}}^{T}({\bm{x}}-\hat{\bm{x}})\right],$ (9)

### v2

@@ 1 Introduction
Denoising is a fundamental problem in classical signal processing and has recently gained renewed attention from the machine learning community. Let ${\bm{x}}=(x_{1},\ldots,x_{n})\in\mathbb{R}_{+}^{n}$ denote a non-negative signal of length $n$, where signal $\bm{x}$} is not observable in many systems. Instead, we observe a noisy version ${\bm{y}}=(y_{1},\ldots,y_{n})$, where the observations are conditionally independent given $\bm{x}$}, and each entry is distributed according to a common conditional distribution:
@@
We assume that the noise mechanism is memoryless (independent across coordinates) and homogeneous (identical across entries). The goal of a denoising algorithm is to estimate $\bm{x}$} from the noisy observations $\bm{y}$}. Given its prevalence in imaging and data acquisition systems, denoising has been a central topic in signal processing for decades. Classical denoising methods rely on explicit structural assumptions about the underlying signal $\bm{x}$}, often hand-crafted by domain experts [wiener1964extrapolation, donoho1994ideal, mallat1999wavelet, donoho2002noising, portilla2003image, elad2006image, roth2009fields, gu2014weighted]. In contrast, recent advances in machine learning have enabled a new class of data-driven denoising algorithms. These methods learn the optimal denoising function from data, leveraging statistical patterns directly from signal and noise distributions.
@@
To mitigate the reliance on paired clean and noisy samples, several self-supervised denoising methods have been developed that learn directly from noisy observations, without access to clean ground truth signals [soltanayev2018training, lehtinen2018noise2noise, krull2019noise2void, batson2019noise2self, kim2021noise2score]. While these approaches alleviate the supervision requirement, they typically depend on access to large collections of noisy data and often yield suboptimal performance compared to methods trained with clean targets. Moreover, the absence of clean supervision necessitates the use of complex neural architectures and training schemes, which can make these methods computationally demanding and difficult to optimize in practice.
@@
In lossy compression, the goal is to represent signals from a target class using discrete encodings with minimal distortion. When applied to noisy data, the intuition is that a lossy compressor—operating at a distortion level matched to the noise—will favor reconstructions close to the original clean signal. While this approach has a strong theoretical foundation [donoho2002kolmogorov, weissman2005empirical], classical compression-based denoisers have shown limited empirical success, particularly for natural image denoising.
In this work, we revisit this idea in light of recent progress in neural compression, where learned encoders and decoders have demonstrated strong rate-distortion performance across a variety of image domains [balle2020nonlinear, yang2023introduction]. Building on this foundation, we propose a zero-shot denoising method that we call the Zero-Shot Neural Compression Denoiser (ZS-NCD). Unlike traditional neural compression models that are trained on large corpora of clean high-resolution images, ZS-NCD learns directly from a single noisy input image. Specifically, we extract overlapping patches from the noisy image, and train a neural compression network on those patches alone—without any clean supervision or prior dataset. Once trained, the denoiser is applied to all patches from the same image, and the final output is obtained by averaging the predictions in overlapping regions. This approach is illustrated in Figure 1.
@@
Figure 2: Zero-shot denoising of Kodim05 with AWGN ($\sigma=25$). Left: PSNR versus training iterations for zero-shot denoisers. Performance of BM3D [dabov2007bm3d] and Restormer [zamir2022restormer] are included as a classical baseline and as a supervised empirical upper bound, respectively. Right: Visual reconstructions with PSNR above each image. Compression-based denoising based on JPEG-2K [taubman2002jpeg2000] achieves inferior performance. Learning-based zero-shot denoisers often struggle with either overfitting or high bias. DIP [dmitry2020dip] and $D$ [heckel2018deep] require early stopping to avoid overfitting. ZS-N2S [batson2019noise2self] and S2S [quan2020self2self] struggle with high-resolution color images, and ZS-N2N [mansour2023zsn2n] often produces noisy outputs with potential overfitting. BM3D tends to oversmooth the denoised image. In contrast, ZS-NCD avoids these issues.
@@ 2 Related work
Supervised learning-based denoisers such as DnCNN [zhang2017beyond] and Restormer [zamir2022restormer] achieve state-of-the-art performance across various noise models, but require large datasets of paired clean and noisy images—often impractical in real-world settings. To avoid clean images, self-supervised methods have been proposed, including Noise2Noise [lehtinen2018noise2noise], Noise2Self [batson2019noise2self], Noise2Void [krull2019noise2void], Noise2Same[xie2020noise2same] and Noise2Score [kim2021noise2score], which only use noisy images for training. However, their reliance on large noisy datasets remains a limitation. Zero-shot denoisers address this by training on a single noisy image. These include (i) untrained networks like DIP [ulyanov2018deep] and Deep Decoder [heckel2018deep], and (ii) single-image adaptations of self-supervised methods, e.g., N2F [lequyer2022fast], ZS-N2N [mansour2023zsn2n], ZS-N2S [batson2019noise2self] and its augmented variant with ensembling, S2S [quan2020self2self]. More recently, DS-N2N [bai2025dual] improves ZS-N2N by further upsampling the downsampled paired noisy images. Pixel2Pixel [ma2025pixel2pixel] boosts the performance by using non-local similarity approach. DIP-based models avoid masking and leverage full-image context, but require early stopping or under-parameterization to avoid overfitting. Self-supervised variants suffer from masking-induced information loss. Hybrid approaches, such as masked pretraining-based method [ma2024masked], uses external datasets for training and perform zero-shot inference, thus falling outside the zero-shot setting studied here.
@@
Learning-based lossy compression, often referred to as neural compression, uses an autoencoder architecture combined with an entropy model to estimate and constrain the bitrate at the bottleneck [balle2017endtoend, theis2017lossy]. These methods have significantly outperformed traditional codecs, particularly in image [balle2017endtoend, balle2018variational, minnen2018joint, zhu2022transformer, liu2023learned] and video [agustsson2020scale, mentzer2022vct, li2024neural] compression. In addition to these standard settings, several works have also explored applying neural compression to noisy data, either by adapting neural compression models for more efficient encoding of noisy images [testolina2021towards, ranjbar2022joint, brummer2023importance, xie2024joint].
@@
Compression-based denoising leverages the insight that structured signals are inherently more compressible than their noisy counterparts. This connection was formalized by Donoho [donoho2002kolmogorov], who introduced the minimum Kolmogorov complexity estimator, and further refined by Weissman et al. [weissman2005empirical], showing that, under certain conditions on both the signal and the noise, optimal lossy compression of a noisy signal-followed by suitable post-processing-can asymptotically achieve optimal denoising performance. Prior to these theoretical developments, early empirical methods such as wavelet-based schemes [saito1994simultaneous, chang1997image, chang2000adaptive] and MDL-inspired heuristics [natarajan1995filtering] had explored this principle. Nevertheless, traditional compression-based denoisers have generally underperformed in high-dimensional settings such as image denoising.
Learning-based joint compression and denoising using neural compression has been explored in recent works [larigauderie2022combining, li2024compress], where the goal is to achieve lower rate in compression. The empirical application of training neural compression for AWGN denoising was also proposed in [zafari2025decompress]. Training the neural compression models in these works requires a dataset of images. In contrast, our proposed ZS-NCD is a two-step denoiser based on neural compression, trained on a single noisy image. It achieves state-of-the-art performance across both AWGN and Poisson noise models. Moreover, we contribute new theoretical results that advance the foundations of compression-based denoising.
@@ 3 Compression-based denoising: Theoretical foundations
Let $\mathcal{Q}\subset\mathbb{R}^{n}$ denote the signal class of interest, such as vectorized natural images of a fixed size. A lossy compression code for $\mathcal{Q}$ is defined by an encoder-decoder pair $(f,g)$, $f:\mathcal{Q}\to\{1,\ldots,2^{R}\}$, and $g:\{1,\ldots,2^{R}\}\to\mathbb{R}^{n}$. The performance of a lossy code is characterized by: i) Rate $R$, indicating the number of distinct codewords; ii) Distortion $\delta$, defined as the worst-case per-symbol mean squared error (MSE) over the signal class:
@@
$\mathcal{C}=\{g(i):i=1,\ldots,2^{R}\}\subset\mathbb{R}^{n}.$
@@
We begin by analyzing the performance of compression-based ML denoising under AWGN. The following result provides a non-asymptotic upper bound on the reconstruction error in terms of the compression rate and distortion. All proofs can be found in Appendix A.
@@
$\frac{1}{\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq\sqrt{\delta}+2\sigma_{z}\sqrt{\frac{(2\ln 2)R}{n}}(1+2\sqrt{\eta}),$ (4)
@@
Let $\mathcal{Q}_{n}$ denote the set of $k$-sparse vectors in $\mathbb{R}^{n}$ satisfying $\|{\bm{x}}\|_{\infty}\leq 1$. Fix a parameter $\eta\in(0,1)$, and suppose ${\bm{y}}={\bm{x}}+{\bm{z}}$ where ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}I_{n})$. Then, there exists a family of compression codes such that, when used with the denoiser defined in (1), with a probability larger than $1-{\rm e}^{-\eta k\log(n/k)}$, the estimate $\hat{\bm{x}}$ satisfies
${1\over\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq\sigma_{z}C\sqrt{{k\over n}\log_{2}({n\over k})+\gamma_{n}}+{1\over\sqrt{n}},$ (5)
where $C=2(1+2\sqrt{\eta})\sqrt{2\ln 2}$ and $\gamma_{n}={k\log_{2}k\over 2n}+{\log_{2}k+k(\log_{2}{\rm e}+1)\over n}$.
Corollary 1 provides a high-probability bound on the normalized error $\frac{1}{\sqrt{n}}\|\hat{{\bm{x}}}-{\bm{x}}\|_{2}$. Squaring both sides and having $(\sqrt{a}+\sqrt{b})^{2}\leq 2(a+b)$, it follows that, with high probability
$\frac{1}{n}\|\hat{{\bm{x}}}-{\bm{x}}\|_{2}^{2}\leq 2\sigma_{z}^{2}C^{2}\Big({k\over n}\log_{2}({n\over k})+\gamma_{n}\Big)+{2\over n}.$
Thus, up to universal constants, the dominant term in the upper bound scales as $\sigma_{z}^{2}{k\over n}\log({n\over k})$. This matches the known minimax rate for estimating $k$-sparse signals in Gaussian noise when $k/n\to 0$; see [johnstone2017gaussian]. Determining whether the residual term $2\over n$} is an artifact of our proof or a real barrier to optimality is an interesting problem. Finally, while the comparison above is high-probability rather than in expectation, one can integrate the tail bound from the proof to obtain an expected-risk bound.
@@
Consider the same setup of lossy compression as in Theorem 1. Assume that for any ${\bm{x}}\in\mathcal{Q}$, $x_{i}\in(x_{\min},x_{\max})$, where $0<x_{\min}<x_{\max}<1$. Assume that $y_{1},\ldots,y_{n}$ are independent with $y_{i}\sim\mathrm{Poisson}(\alpha x_{i})$. Let $\hat{\bm{x}}$ denote the solution of (2). Let $C_{1}={x_{\max}^{5}/(x_{\min}^{2})}$ and $C_{2}={x_{\max}^{2}\over x_{\min}^{3}}\beta\sqrt{({4\over\ln 2})}(\sqrt{1+\eta}+\sqrt{\eta})$. Then, with a probability larger than $1-2^{-\eta R+2}$,
@@
Consider the same setup as in Theorem 2. Let $\hat{\bm{x}}$ denote the solution of (3). Let $C=4\sqrt{\ln 2}(\sqrt{1+\eta}+\sqrt{\eta}+1)$. Then, with a probability larger than $1-2^{-\eta R+2}$,
${1\over n}\|{\bm{x}}-\hat{\bm{x}}\|_{2}^{2}$ $\leq\delta+C\sqrt{R\over n\alpha}.$ (7)
@@
Theorems 2 and  3 show that, in the case of Poisson noise, minimizing either the ML loss function or the computationally efficient MSE loss function can recover the signal. This result is also consistent with our simulations reported later in Section 5.
@@ 4 Zero-shot compression-based denoiser
Let $\mathcal{P}_{(i,j)}:\mathbb{R}^{h\times w}\to\mathbb{R}^{k\times k}$ denote the patch extraction operator, which returns a $k\times k$ patch whose top-left corner is at pixel $(i,j)$. Let $f_{\bm{\theta}_{1}}$ and $g_{\bm{\theta}_{2}}$ denote the encoder and decoder networks, parameterized by weights $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, respectively. Define $\mathcal{I}$ as the set of all coordinates $(i,j)\in\{1,\ldots,h-k+1\}\times\{1,\ldots,w-k+1\}$ from which a valid $k\times k$ patch can be extracted.
@@
[…]
@@
1:Initialize: $\lambda^{(0)}>0$, $\texttt{tol}>0$, $\zeta\in(0,1)$
2:for $k=0,1,\dotsc,K_{\max}-1$ do
3:  Estimate $\hat{\bm{x}}^{(k)}$ from (9).
4:  Compute residual $r^{(k)}=\|\hat{\bm{x}}^{(k)}-\bm{y}\|^{2}$.
5:  Compute scale factor $\beta=(r^{(k)}-n\sigma_{z}^{2})/n\sigma_{z}^{2}$.
6:  if $|\beta|\leq\texttt{tol}$ then
7:   break
8:  else
9:
$\lambda^{(k+1)}:=\begin{cases}\lambda^{(k)}/(1+\zeta|\beta|),&\;\beta>0\\ \lambda^{(k)}(1+\zeta|\beta|),&\;\text{otherwise}\end{cases}$
10:  end if
11:end for
12:Return: $\lambda^{*}=\lambda^{(k)}$
@@
Setting the hyperparameter $\bm{\lambda}$
@@
The ZS-NCD objective in (8) includes a hyperparameter, $\lambda$, which balances reconstruction fidelity and compressibility. Interpreted through the lens of lossy compression, varying $\lambda$ allows the model to explore different rate-distortion trade-offs. However, in the context of denoising, our goal is not compression but accurate signal recovery from the noisy observation $\bm{y}$}. This raises the central question: how should $\lambda$ be selected to optimize denoising performance? In the following, we explain our approach for setting $\lambda$ under both AWGN model and Poisson noise model.
@@
Case I: Gaussian noise. Let $\hat{\bm{x}}$ denote the output of the ZS-NCD denoiser, and consider the AWGN model ${\bm{y}}={\bm{x}}+{\bm{z}}$ with ${\bm{z}}\sim\mathcal{N}(0,\sigma_{z}^{2}I_{n})$. Then,
[…]